# Stong order $1$ adaptive approximation of jump-diffusion SDEs with discontinuous drift

### v1

@@ Abstract
We present an adaptive approximation scheme for jump-diffusion SDEs with discontinuous drift and (possibly) degenerate diffusion. This transformation-based doubly-adaptive quasi-Milstein scheme is the first scheme that has strong convergence rate $1$ in $L^{p}$ for $p\in[1,\infty)$ with respect to the average computational cost for these SDEs. To obtain our result, we prove that under slightly stronger assumptions which are still weaker than those in existing literature, a related doubly-adaptive quasi-Milstein scheme has convergence order $1$. This scheme is doubly-adaptive in the sense that it is jump-adapted, i.e. all jump times of the Poisson noise are grid points, and it includes an adaptive stepsize strategy to account for the discontinuities of the drift.
@@ 1 Introduction
where $\xi\in{\mathbb{R}}$, $\mu,\sigma,\rho\colon{\mathbb{R}}\to{\mathbb{R}}$ are measurable functions, $T\in(0,\infty)$, $W=(W_{t})_{t\geq 0}$ is a standard Brownian motion, and $N=(N_{t})_{t\geq 0}$ is a homogeneous Poisson process with intensity $\lambda\in(0,\infty)$ on the probability space $(\Omega,{\cal F},{\mathbb{P}})$. Further, the filtration $\mathbb{F}=(\mathbb{F}_{t})_{t\geq 0}$ is the augmented natural filtration of $N$ and $W$, i.e. for all $t\geq 0$, $\mathbb{F}_{t}=\sigma(\{(W_{s},N_{s})\colon s\leq t\}\cup\mathcal{N})$, where $\mathcal{N}=\{A\in{\cal F}\colon{\mathbb{P}}(A)=0\}$. Then $(\Omega,{\cal F},\mathbb{F},{\mathbb{P}})$ is a filtered probability space satisfying the usual conditions.
We study the numerical approximation of these SDEs in the case of a discontinuous drift. Under our assumptions existence and uniqueness in the strong sense is settled by [54, 55]. Such SDEs are, for example, used for modelling energy prices or control actions on energy markets, cf. [63, 49, 1, 61, 62]. We construct an adaptive approximation scheme that has convergence rate $1$ for jump-diffusion SDEs with discontinuous drift. The scheme is a transformation-based doubly-adaptive quasi-Milstein scheme, which combines for the first time two different types of adaptivity. It is jump-adapted, i.e. all the jump times of the Poisson process are added to the grid, and it is adaptive with respect to the discontinuity points of the drift, i.e. the closer the approximated solution is to a discontinuity point, the smaller the size of the next time step. This is the first approximation scheme for jump-diffusion SDEs with discontinuous drift that has convergence rate $1$ in $L^{p}$ for $p\in[1,\infty)$ and hence recovers the optimal rate for non-adaptive as well as adaptive schemes for a large class of SDEs without jump noise and with Lipschitz-continuous coefficients [35, 36].
@@
The transformation-based doubly-adaptive quasi-Milstein scheme defined in this article is inspired by some ideas from the literature on the construction of algorithms. Transformation-based schemes have been introduced in [30, 31] for SDEs with discontinuous drift without a jump noise. The transformation is defined in such a way that the drift coefficient of the transformed SDE gains additional regularity, but is still less regular than in previously existing literature. This technique inspired many of the subsequent contributions. A transformation-based quasi-Milstein scheme is proven to have convergence order $3/4$ in $L^{p}$ in [38] for SDEs with discontinuous drift without a jump noise. Schemes with adaptive step-size function originate in [44], where an adaptive Euler–Maruyama scheme has been studied for the approximation of SDEs with discontinuos drift without a jump noise. A similar idea has been used in [65] to construct a transformation-based adaptive quasi-Milstein scheme with convergence rate $1$ in $L^{p}$ for SDEs with discontinuous drift without jump noise. The schemes and results in [38, 65] cannot be extended straight forward to jump-diffusion SDEs. Instead jump-adapted approximation schemes are needed, which are for example studied in [49, 28, 57]. In [57] the transformation-based jump-adapted quasi-Milstein scheme has been studied for jump-diffusion SDEs but the best rate that could be achieved by this scheme is $3/4$.
In this paper, we make use of these ideas to define the first doubly-adaptive approximation scheme and prove that it converges with strong rate $1$. The proof is performed in two steps. First, under slightly stronger assumptions, which are still weaker than in the existing literature, we define a doubly-adapted quasi-Milstein scheme and prove convergence rate $1$. Our assumptions require a careful derivation of occupation time estimates. Then we use a transformation to construct the first rate $1$ approximation scheme for jump-diffusion SDEs with discontinuous drift, the so-called transformation-based doubly-adaptive quasi-Milstein scheme, and prove its convergence rate.
@@
In Section 2 we define the setting. In Section 3 we define the doubly-adaptive quasi-Milstein scheme, prove its strong convergence rate and estimate the computational cost. In Section 4 we define based on the previous section the transformation-based doubly-adaptive quasi-Milstein scheme and prove the main result of this article.
@@ 2 Setting
In the following, we denote the compensated Poisson process by $\tilde{N}=(\tilde{N}_{t})_{t\geq 0}$ with $\tilde{N}_{t}=N_{t}-\lambda t$ for all $t\in[0,\infty)$. Note that the compensated Poisson process is a square-integrable $\mathbb{F}$-martingale. Further, let $L_{f}$ be the Lipschitz constant of a generic Lipschitz continuous function $f$, and $\operatorname{Id}$ the identity mapping. We recall the definition of piecewise Lipschitz:
@@
Assume that there exist $m\in{\mathbb{N}}$ and $-\infty=\zeta_{0}<\zeta_{1}<\ldots<\zeta_{m}<\zeta_{m+1}=\infty$ such that the coefficients of SDE (1.1) satisfy:
@@
    $\mu$ and $\sigma$ are differentiable with Lipschitz continuous derivatives on $(\zeta_{i-1},\zeta_{i})$ for all $i\in\{1,\ldots,m\}$.
Remark 2.2.
Under Assumption 2.1 it is well-known that $\mu$, $\sigma$, and $\rho$ satisfy a linear growth condition, i.e. for $f\in\{\mu,\sigma,\rho\}$ there exist constants $c_{f}\in(0,\infty)$ such that $|f(x)|\leq c_{f}(1+|x|)$ and that there exists a unique strong solution to SDE (1.1), e.g. [54, Lemma 2].
@@ 3 Convergence of the doubly-adaptive quasi-Milstein scheme
    there exists $\widetilde{m}\in{\mathbb{N}}$ and $-\infty=\zeta_{0}<\zeta_{1}<\ldots<\zeta_{\widetilde{m}}<\zeta_{\widetilde{m}+1}=\infty$ such that $\widetilde{\mu}$ and $\widetilde{\sigma}$ are differentiable with Lipschitz continuous derivatives on $(\zeta_{k-1},\zeta_{k})$ for $k\in\{1,\ldots,\widetilde{m}+1\}$;
@@
$d_{f}(x)=\left\{\begin{array}[]{ll}f^{\prime}(x)&f\text{ differentiable at }x\\ 0&\,\textrm{else.}\\ \end{array}\right.$ (3.2)
@@
Here the sequence of grid points $(\tau_{n})_{n\in{\mathbb{N}}}$ is chosen in a two-fold adaptive way. Namely it contains all jump points of the Poisson process and it takes into account how far the last approximated value of the solution is from the discontinuities (due to the stepsize function $h^{\delta}$). Note that the following integral notation is equivalent to (3.6), (3.8), (3.9), (3.10),
@@ 3.1 Preparatory lemmas
where $\mathcal{N}$ is the set of all nullsets of ${\cal F}$. $W$ is a Brownian motion with respect to the filtration $(\widetilde{\mathbb{F}}_{t})_{t\geq 0}$ as $W$ and $N$ are independent. Moreover, $Z^{(M)}$ and $Z$ are $\widetilde{\mathbb{F}}$-adapted càdlàg processes, since $\mathbb{F}_{t}\subset\widetilde{\mathbb{F}}_{t}$ for all $t\in[0,\infty)$.
@@
Let $\delta\in(0,\delta_{0}]$ then the sequence $(\tau_{n})_{n\in{\mathbb{N}}}$ satisfies for all $n\in{\mathbb{N}}$,
@@
By induction, we prove (i) and (ii). Note that $\tau_{0}\equiv 0$ is a $\mathbb{F}$-stopping time and $Z^{(\delta)}_{\tau_{0}}\equiv\widetilde{\xi}$ is $\mathbb{F}_{0}=\mathbb{F}_{\tau_{0}}$-measurable. Further, if $\tau_{n}$ is a stopping time and $Z^{(\delta)}_{\tau_{n}}$ is $\mathbb{F}_{\tau_{n}}$-measurable, then $\tau_{n}+h^{\delta}(Z^{(\delta)}_{\tau_{n}})>\tau_{n}$ is obviously $\mathbb{F}_{\tau_{n}}$-measurable and hence a $\mathbb{F}$-stopping time, cf. [27, Exercise 1.2.14]. This implies that $\tau_{n+1}=\big{(}\tau_{n}+h^{\delta}(Z^{(\delta)}_{\tau_{n}})\big{)}\wedge\min\{\nu_{i}:i\in{\mathbb{N}},\nu_{i}>\tau_{n}\}\wedge T$ is a $\mathbb{F}$-stopping time. Hence, by construction $Z^{(\delta)}_{\tau_{n+1}}$ is $\mathbb{F}_{\tau_{n+1}}$-measurable. Item (iii) is implied by [60, Theorem 40.10] considering the two-dimensional Lévy process $(N,W)$ and (iv) is proven in [24, Theorem 11.11]. ∎
@@
Note that $\underline{t}$ is a random time, but no stopping time. An introduction to random times can, for example, be found in [27, p. 5]. Further we define for all $s,t\in[0,\infty)$ the random variable $W^{\underline{t}}_{s}=W_{\underline{t}+s}-W_{\underline{t}}$ and the stochastic process $W^{\underline{t}}=(W^{\underline{t}}_{s})_{s\geq 0}$. Note that $W^{\underline{t}}$ is ${\mathbb{P}}$-a.s. continuous.
@@
For all $A\in\mathcal{B}(C([0,\infty),{\mathbb{R}}))$ we have by Lemma 3.3 (ii), (iii), and the fact that $\tau_{n+1}$ is $\widetilde{F}_{\tau_{n}}$-measurable that
@@
Let Assumption 3.1 hold. Then for all $\delta\in(0,\delta_{0}]$, $M=\Big{\lceil}\frac{T}{\delta^{2}\log^{4}(\delta^{-1})}\Big{\rceil}$ and $p\in[2,\infty)$ there exists a constant $c_{\delta}\in(0,\infty)$ such that
$\int_{0}^{T}{\mathbb{E}}\big{[}1+\big{|}Z^{(\delta)}_{\underline{t}}\big{|}^{p}\big{]}\mathop{}\!\mathrm{d}t+{\mathbb{E}}\Big{[}N_{T}^{p-1}\sum_{n=1}^{N_{T}+M}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}-}\big{|}^{p}\big{)}\Big{]}\leq c_{\delta}.$ (3.23)

### v2

# Strong order $1$ adaptive approximation of jump-diffusion SDEs with discontinuous drift

@@ Abstract
We present an adaptive approximation scheme for jump-diffusion SDEs with discontinuous drift and (possibly) degenerate diffusion. This transformation-based doubly-adaptive quasi-Milstein scheme is the first scheme that has strong convergence rate $1$ in terms of the number of evaluations of the driving noise processes in $L^{p}$ for $p\in[1,\infty)$. To obtain our result, we prove that under slightly stronger assumptions, which are still weaker than those in the existing literature, a related doubly-adaptive quasi-Milstein scheme has convergence order $1$. This scheme is doubly-adaptive in the sense that it is jump-adapted, i.e. all jump times of the Poisson noise are grid points, and it includes an adaptive step-size strategy to account for the discontinuities of the drift.
@@ 1 Introduction
where $\xi\in{\mathbb{R}}$, $\mu,\sigma,\rho\colon{\mathbb{R}}\to{\mathbb{R}}$ are measurable functions, $T\in(0,\infty)$, $W=(W_{t})_{t\geq 0}$ is a standard Brownian motion, and $N=(N_{t})_{t\geq 0}$ is a homogeneous Poisson process with intensity $\lambda\in(0,\infty)$ on the complete probability space $(\Omega,{\cal F},{\mathbb{P}})$. Further, the filtration $\mathbb{F}=(\mathbb{F}_{t})_{t\geq 0}$ is the augmented natural filtration of $N$ and $W$, i.e. for all $t\geq 0$, $\mathbb{F}_{t}=\sigma(\{(W_{s},N_{s})\colon s\leq t\}\cup\mathcal{N})$, where $\mathcal{N}=\{A\in{\cal F}\colon{\mathbb{P}}(A)=0\}$. Then $(\Omega,{\cal F},\mathbb{F},{\mathbb{P}})$ is a filtered probability space satisfying the usual conditions.
We study the numerical approximation of these SDEs in the case of a discontinuous drift. To be more precise, we assume that the drift coefficient is piecewise Lipschitz, the diffusion and jump coefficients are Lipschitz continuous, the diffusion is non-zero at the discontinuities of the drift, and the drift and diffusion coefficients are differentiable with Lipschitz continuous derivatives in between neighbouring discontinuities. Under our assumptions existence and uniqueness in the strong sense is settled by [54, 55]. Such SDEs are, for example, used for modelling energy prices or control actions on energy markets, cf. [63, 49, 1, 61, 62]. We construct an adaptive approximation scheme that has convergence rate $1$ in terms of the number of evaluations of the driving noise processes for jump-diffusion SDEs with discontinuous drift, the so-called transformation-based doubly-adaptive quasi-Milstein scheme. The main novelty is that for the first time, we combine two different types of adaptivity in this scheme. It is jump-adapted, i.e. all the jump times of the Poisson process are added to the grid, and it is adaptive with respect to the discontinuity points of the drift, i.e. the closer the approximated solution is to a discontinuity point, the smaller the size of the next time step. This is the first approximation scheme for jump-diffusion SDEs with discontinuous drift that has convergence rate $1$ in terms of the number of evaluations of the driving noise processes in $L^{p}$ for $p\in[1,\infty)$ and hence recovers the optimal rate for non-adaptive as well as adaptive schemes for a large class of SDEs without jump noise and with Lipschitz-continuous coefficients [35, 36].
@@
The transformation-based doubly-adaptive quasi-Milstein scheme defined in this article is inspired by some ideas from the literature on the construction of algorithms. Transformation-based schemes have been introduced in [30, 31] for SDEs with discontinuous drift without a jump noise. The transformation is defined in such a way that the drift coefficient of the transformed SDE gains additional regularity. This technique inspired many of the subsequent contributions. A transformation-based quasi-Milstein scheme is proven to have convergence order $3/4$ in $L^{p}$ in [38] for SDEs with discontinuous drift without a jump noise. Schemes with adaptive step-size function originate in [44], where an adaptive Euler–Maruyama scheme has been studied for the approximation of SDEs with discontinuos drift without a jump noise. A similar idea has been used in [65] to construct a transformation-based adaptive quasi-Milstein scheme with convergence rate $1$ in terms of the number of evaluations of the driving Brownian motion in $L^{p}$ for SDEs with discontinuous drift without jump noise. The schemes and results in [38, 65] cannot be extended straightforwardly to jump-diffusion SDEs. Instead jump-adapted approximation schemes are needed, which are for example studied in [49, 28, 57]. In [57] the transformation-based jump-adapted quasi-Milstein scheme has been studied for jump-diffusion SDEs but the best rate that could be achieved by this scheme is $3/4$.
In this paper, we combine the approximation schemes from [65] and [57], which are based on ideas introduced in [44] resp. [49], to define the first doubly-adaptive approximation scheme and prove that it converges with strong rate $1$ in terms of the number of evaluations of the driving noise processes. For this a transformation $G\colon{\mathbb{R}}\to{\mathbb{R}}$ is applied to $X$ to obtain an SDE for $Z=G(X)$, which satisfies the assumptions made on the coefficients of $X$ and additionally has a continuous drift. This implies that the drift of the transformed SDE is Lipschitz continuous. For the approximation of the solution $Z$ we define the same step-size function $h^{\delta}\colon{\mathbb{R}}\to{\mathbb{R}}$ as in [65], which has value $\delta$ away from the discontinuities of the drift and smaller values closer to the discontinuities of the drift. Given the value of the approximation $Z^{(\delta)}$ at some grid point, we determine the next grid point by adding the minimum of the step-size function evaluated at this point and the time until the next jump of the Poisson process. Then we compute the left limit of $Z^{(\delta)}$ at the next grid point via a quasi-Milstein scheme. In case this grid point is a jump time of the Poisson process, we then add the corresponding jump term; for details see (3.6)–(3.9).
For this scheme we prove in Theorem 3.16 that under our assumptions for all $p\in[1,\infty)$ there exists a constant $c\in(0,\infty)$ such that for small enough values of $\delta$,
${\mathbb{E}}\Big[\sup_{t\in[0,T]}\big|Z(t)-Z^{(\delta)}(t)\big|^{p}\Big]^{\frac{1}{p}}\leq c\delta.$
To ensure that the cost of our adaptive algorithm stays proportional to $\delta$, a cost analysis is necessary. For this, we denote by $n(Z^{(\delta)}_{T})$ the maximum number of time steps needed to reach $T$, which is proportional to the computational cost of the numerical approximation $Z^{(\delta)}$. In Theorem 3.17 we prove that there exists a constant $\widetilde{c}\in(0,\infty)$ such that for small enough values of $\delta$,
${\mathbb{E}}[n(Z^{(\delta)}_{T})]\leq c(\delta^{-1}+{\mathbb{E}}[N_{T}]).$
Using these results and analytical properties of the transformation $G$ and its inverse $G^{-1}$ we prove in Theorem 4.3 under our assumptions on the coefficients of SDE (1.1) that there exist $C,\widetilde{C}\in(0,\infty)$ such that for small enough values of $\delta$,
${\mathbb{E}}\Big[\sup_{t\in[0,T]}|X_{t}-G^{-1}(Z^{(\delta)}_{t})|^{p}\Big]^{\frac{1}{p}}\leq C\delta,\qquad{\mathbb{E}}\Big[n\big(G^{-1}(Z^{(\delta)}_{T})\big)\Big]\leq\widetilde{C}(\delta^{-1}+{\mathbb{E}}[N_{T}]).$ (1.2)
Note that in case $\rho\equiv 0$ our scheme coincides with the scheme considered in [65] and our proofs follow the general strategy presented in [65]. However, adding the driving Poisson noise and the corresponding jump coefficient requires important changes in the definition of the approximation scheme compared to [65]. Similar to [57] a jump-adapted grid needs to be included in the scheme. In combination with the adaptive step-size function, this two-fold adaptivity is the main novelty of the scheme. Together with the jump noise itself, they present challenges for the theoretical proofs. To meet these challenges, we introduce a series representation of the coefficients of the SDE and the approximation scheme, we condition on either the Poisson noise having a certain value or on additionally introduced filtrations, and we carefully estimate the newly appearing terms. The big advantage of using a jump-adapted scheme is that the approximation scheme jumps at exactly the same times as the solution itself. This allows for improved estimates on the jump terms of the approximation scheme, which are crucial to obtain strong convergence rate $1$ in terms of the number of evaluations of the driving noise processes in $L^{p}$ for $p\in[1,\infty)$.
@@
The paper is structured as follows: In Section 2 we introduce the setting. In Section 3 we define the doubly-adaptive quasi-Milstein scheme, prove its strong convergence rate and estimate the computational cost. In Section 4 we define based on the previous section the transformation-based doubly-adaptive quasi-Milstein scheme and prove the main result of this article.
@@ 2 Setting
In the following, we denote the compensated Poisson process by $\tilde{N}=(\tilde{N}_{t})_{t\geq 0}$ with $\tilde{N}_{t}=N_{t}-\lambda t$ for all $t\in[0,\infty)$. Note that the compensated Poisson process is a square-integrable $\mathbb{F}$-martingale. Further, let $L_{f}$ be the Lipschitz constant of a generic Lipschitz continuous function $f$, and $\operatorname{Id}$ the identity mapping. Moreover, for two measurable spaces $(A_{1},\mathcal{A}_{1})$ and $(A_{2},\mathcal{A}_{2})$, for measurable functions $Y_{1}\colon(\Omega,\mathcal{F})\to(A_{1},\mathcal{A}_{1})$, $Y_{2}\colon(\Omega,\mathcal{F})\to(A_{2},\mathcal{A}_{2})$, and $f\colon A_{1}\times A_{2}\to{\mathbb{R}}$, and $\mathcal{G}$ a sub-$\sigma$-algebra of $\mathcal{F}$, we denote by ${\mathbb{E}}[f(Y_{1},y)|\mathcal{G}]|_{y=Y_{2}}$ the random variable $g(Y_{2})$, where $g\colon A_{2}\to{\mathbb{R}}$ is defined by $g(y)={\mathbb{E}}[f(Y_{1},y)|\mathcal{G}]$ for all $y\in A_{2}$. In particular, this notation is used in case that $\mathcal{G}=\{\emptyset,\Omega\}$, then conditioning is omitted and we write ${\mathbb{E}}[f(Y_{1},y)]|_{y=Y_{2}}$. We recall the definition of piecewise Lipschitz:
@@
Assume that there exist $m\in{\mathbb{N}}$ and $\infty=\zeta_{0}<\zeta_{1}<\ldots<\zeta_{m}<\zeta_{m+1}=\infty$ such that the coefficients of SDE (1.1) satisfy:
@@
    $\mu$ and $\sigma$ are differentiable with Lipschitz continuous derivatives on $(\zeta_{i-1},\zeta_{i})$ for all $i\in\{1,\ldots,m+1\}$.
Remark $2.$
Under Assumption 2.1 it is well-known that $\mu$, $\sigma$, and $\rho$ satisfy a linear growth condition, i.e. for $f\in\{\mu,\sigma,\rho\}$ there exists a constant $c_{f}\in(0,\infty)$ such that $|f(x)|\leq c_{f}(1+|x|)$ and that there exists a unique strong solution to SDE (1.1), e.g. [54, Lemma 2].
@@ 3 Convergence of the doubly-adaptive quasi-Milstein scheme
    there exists $\widetilde{m}\in{\mathbb{N}}$ and $\infty=\zeta_{0}<\zeta_{1}<\ldots<\zeta_{\widetilde{m}}<\zeta_{\widetilde{m}+1}=\infty$ such that $\widetilde{\mu}$ and $\widetilde{\sigma}$ are differentiable with Lipschitz continuous derivatives on $(\zeta_{k-1},\zeta_{k})$ for $k\in\{1,\ldots,\widetilde{m}+1\}$;
@@
[…]
@@
Here the sequence of grid points $(\tau_{n})_{n\in{\mathbb{N}}}$ is chosen in a two-fold adaptive way. Namely it contains all jump points of the Poisson process and it takes into account how far the last approximated value of the solution is from the discontinuities of $\mu$ (due to the step-size function $h^{\delta}$). Note that the following integral notation is equivalent to (3.6), (3.8), (3.9), (3.10),
@@ 3.1 Preparatory lemmas
[…]
@@
Let $\delta\in(0,\delta_{0}]$. Then the sequence $(\tau_{n})_{n\in{\mathbb{N}}}$ satisfies for all $n\in{\mathbb{N}}$,
@@
By induction, we prove (i) and (ii). Note that $\tau_{0}\equiv 0$ is an $\mathbb{F}$-stopping time and $Z^{(\delta)}_{\tau_{0}}\equiv\widetilde{\xi}$ is $\mathbb{F}_{0}=\mathbb{F}_{\tau_{0}}$-measurable. Further, if $\tau_{n}$ is a stopping time and $Z^{(\delta)}_{\tau_{n}}$ is $\mathbb{F}_{\tau_{n}}$-measurable, then $\tau_{n}+h^{\delta}(Z^{(\delta)}_{\tau_{n}})>\tau_{n}$ is obviously $\mathbb{F}_{\tau_{n}}$-measurable and hence an $\mathbb{F}$-stopping time, cf. [27, Exercise 1.2.14]. This implies that $\tau_{n+1}=\big(\tau_{n}+h^{\delta}(Z^{(\delta)}_{\tau_{n}})\big)\wedge\min\{\nu_{i}:i\in{\mathbb{N}},\nu_{i}>\tau_{n}\}\wedge T$ is a $\mathbb{F}$-stopping time. Hence, by construction $Z^{(\delta)}_{\tau_{n+1}}$ is $\mathbb{F}_{\tau_{n+1}}$-measurable. Item (iii) is implied by [60, Theorem 40.10] considering the two-dimensional Lévy process $(N,W)$ and (iv) is proven in [24, Theorem 11.11]. ∎
@@
Note that $\underline{t}$ is a random time, but not a stopping time. An introduction to random times can, for example, be found in [27, p. 5]. Further we define for all $s,t\in[0,\infty)$ the random variable $W^{\underline{t}}_{s}=W_{\underline{t}+s}-W_{\underline{t}}$ and the stochastic process $W^{\underline{t}}=(W^{\underline{t}}_{s})_{s\geq 0}$. Note that $W^{\underline{t}}$ is $\mathbb{P}$}-a.s. continuous.
@@
For all $A\in\mathcal{B}(C([0,\infty),{\mathbb{R}}))$ we have by Lemma 3.3 (ii), (iii), and the fact that $\tau_{n+1}$ is $\widetilde{\mathbb{F}}_{\tau_{n}}$-measurable that
@@
Let Assumption 3.1 hold. Then for all $\delta\in(0,\delta_{0}]$ and $p\in[2,\infty)$ there exists a constant $c_{\delta}\in(0,\infty)$ such that
$\int_{0}^{T}{\mathbb{E}}\big[1+\big|Z^{(\delta)}_{\underline{t}}\big|^{p}\big]\mathop{}\!\mathrm{d}t+{\mathbb{E}}\Big[N_{T}^{p-1}\sum_{n=1}^{N_{T}+M}\big(1+\big|Z^{(\delta)}_{\tau_{n}-}\big|^{p}\big)\Big]\leq c_{\delta},$ (3.23)
where $M=\Big\lceil\frac{T}{\delta^{2}\log^{4}(\delta^{-1})}\Big\rceil$.